OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 8 inches.

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

What is the volume of water in the vase?

OpenStudy (anonymous):

what is the total volume of marbles?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

total filled volume of vase is \[V _{tot}=\frac{\pi D^2}{4} H=\frac{16\pi}{4}*8=32\pi\]

OpenStudy (anonymous):

\[V_{marbles}=6*\frac{4}{3} \pi r^3 \ \ \ \ r=d/2=2/2=1 \\ then \ \ \ V_{marbles}=6*\frac{4}{3} \pi =8\pi\]

OpenStudy (anonymous):

\[V_{water}=V_{tot}V_{marbles}=32\pi-8\pi=24\pi\]

OpenStudy (anonymous):

you just made this really confusing i thought it was πr2h

OpenStudy (anonymous):

πr2h is total volume not the volume of water

OpenStudy (anonymous):

***V(water)=V(Total)-V(marbles)